Which mixed number is equivalent to the improper fraction?
41/7

Answer by stanbon(75887)

Which mixed number is equivalent to the improper fraction?
41/7 = (35/7)+(6/7) = 5 6/7
===============================
Cheers,
Stan H.
==========